OpenStudy (brucebaner):

What is the value of the expression? 4^3 • 4^−5 A. 4^−15 B. 4^−2 C. 4^2 D. 4^8

OpenStudy (brucebaner):

@Michele_Laino

OpenStudy (alexandervonhumboldt2):

4^(3-5)

OpenStudy (alexandervonhumboldt2):

4^-2

OpenStudy (alexandervonhumboldt2):

a^b*a^c=a(b+c) (a^b)^c=a^bc

OpenStudy (michele_laino):

\[4^{3}*4^{-5}=4^{(3-5)}=4^{-2}=\frac{ 1 }{ 4^{2} }=\frac{ 1 }{ 16 }\]

OpenStudy (alexandervonhumboldt2):

4^-2

OpenStudy (brucebaner):

so what letter is it

OpenStudy (alexandervonhumboldt2):

B

OpenStudy (michele_laino):

please, @brucebaner are all my step clear for you?

OpenStudy (brucebaner):

what about you @saifoo.khan

OpenStudy (saifoo.khan):

I'd go with @Michele_Laino .

OpenStudy (brucebaner):

so its B

OpenStudy (michele_laino):

sorry please, I want to know if my reasoning is clear for you!

OpenStudy (brucebaner):

yes

OpenStudy (michele_laino):

ok! then what properties of power I applied before?

OpenStudy (brucebaner):

yes i your reasoning is clear for me

OpenStudy (michele_laino):

ok! please solve this little test and after that I will give you your answer. evaluate this, please: \[\frac{ 5^{-2} }{ 5^{-3} }\]

OpenStudy (brucebaner):

5^-5

OpenStudy (michele_laino):

please, retry!

OpenStudy (michele_laino):

please note that: \[\frac{ 1 }{ 2^{-9}}=2^{9}\]

OpenStudy (michele_laino):

that's right! your answer is B.

OpenStudy (brucebaner):

thx

OpenStudy (michele_laino):

@brucebaner sorry where is your correct answer?